OpenStudy (anonymous):

Determine if statement (3) follows from statements (1) and (2). if it does, state the law that explains why. If it does not, write "invalid". (1) parallel lines do not intersect (2) lines m and n are parallel (3) lines m and n do not intersect

OpenStudy (anonymous):

i think its valid by the law of syllogism. is that right or...?

OpenStudy (perl):

the law of syllogism is a little different. $$ \Large{ if~p~then~ q \\ if~q~then~ r \over if~ p ~then ~r } $$

OpenStudy (anonymous):

so then it's invalid?

OpenStudy (perl):

it is valid, but using a different rule

OpenStudy (anonymous):

detachment?

OpenStudy (perl):

yes

OpenStudy (anonymous):

thank you!

OpenStudy (perl):

law of detachment: $$ \Large{ if~p,~then~ q \\ ~~p \over q } $$

OpenStudy (anonymous):

think you can help me with 1 more question?

OpenStudy (perl):

ok

OpenStudy (anonymous):

Write a two-column proof for the following: given:∠1 and ∠2 form a linear pair and m∠1=2(m∠2) prove:m∠2=60

OpenStudy (anonymous):

sorry, given: angle1 and angle2 form a linear pair and angle1=2(angle2) prove:angle2=60

OpenStudy (anonymous):

i know that a linear pair adds up to 180 degrees, so angles 1 and 2 are supplementary, so 180-120=60, but idk how to write this in proof form. and i have to state the properties/laws that go with each step

OpenStudy (perl):

can you upload an image of the assignment, it is hard to follow the directions

OpenStudy (anonymous):

it's question 9

OpenStudy (perl):

thanks

OpenStudy (perl):

<1,<2 linear pair (given) <1 +<2 = 180 (linear pair are supplementary) (2*<2) + <2 = 180 (by substitution) 3*<2 = 180 (by algebra) <2 = 180 / 3 (by algebra)

OpenStudy (anonymous):

what do you mean by algebra?

OpenStudy (perl):

collect like terms

OpenStudy (anonymous):

ok ty!

OpenStudy (perl):

do you see the substitution part, because in the directions <1 = 2* <2

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

but where did you get the 3*<2 part

OpenStudy (perl):

<1,<2 linear pair (given) m<1 +m<2 = 180 (linear pair are supplementary) 2*m<2 + m<2 = 180 (substitution) 3*m<2 = 180 (addition, collect like terms ) m<2 = 180 / 3 (equals divided by equals is equal)

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

ok i got it, thx again!